An electron is to be accelerated from 2.0 x 106m/s to 7.5 x 106m/s. Through what potential difference must the electron pass to accomplish this?


1
Expert's answer
2020-08-18T11:17:50-0400

As per the question,

Let the initial velocity be v"_1" =2"\times10^6" m/s

and the final velocity be v"_2=7.5\times10^6m/s"

As we know

Mass of electron, m="9.1\times10^{-31}" kg

charge of electron , q="1.6\times10^{-19}" c

Change in kinetic energy= Charge of electron"\times" potential applied

"\frac{mv_2^2}{2}-\frac{mv_1^2}{2}=q\times V"

"\frac{m(v_2^2-v_1^2)}{2}= q\times V"

"\frac{9.1\times10^{-31}(7.5^2-2^2)\times10^{12}}{2}=1.6\times10^{-19}\times V"

On solving the above equation we get,

potential V=148.58volts.